At room temperature (20°C) milk turns sour in about 64 h. In a refrigerator at 3°C
milk can stored three times as long before it sours. How long should it take milk to
sour at 40°C.

Grade:12th pass

1 Answers

Arun
6 years ago
Dear student,
 
Given 
K (293)/K(276) = 3
T2 = 293 K, T1 = 276 K
 
2.303 log (K2/K1) = Ea(T2 -T1)/R* (T1)(T2)
On solving
Ea = 10.454 Kcal
 
Also
2.303 log(K3/K2) = Ea (T3-T2)/2* (T3)(T2)
Where T3 = 313 K
 
Now on solving
K3 = 3.12
 
K3/K2 = t2/t3
Where t2 = 64 Hr., We have to find t3
Now
t3 = 64*1/3.12 = 20.5 hr
 
Regards
Arun (askIITians forum expert)